40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17056     Entry time: Wed Aug 3 16:00:51 2022     In reply to: 17046     Reply to this: 17067   17208   17277
Author: yuta 
Type: Update 
Category: BHD 
Subject: BHD fringe aligned with reduced LO and AS beam clipping 

Last week, we could find an alignment which realizes LO beam and AS beam both unclipped, but it was not consistent with an alignment which realize BHD fringe (40m/17046).
Today, we tweaked the alignment of SR2, AS1, AS4 to have BHD fringe with reduced LO and AS beam clipping.
AS beams on AP table and BHD both still look clipped, but much better now.
Ideally, SR2 and AS1 will unclip AS beam, and LO1, LO2, AS4 would make BHD fringe, but it is hard right now since LO beam seem to have little room and LO2 have little actuation range.
BHD optics on ITMY table, including camera, and AS55/ASDC were realigned after the aglinment work (Note that DCPD_A path have a pick-off for camera path, and this pick-off mirror have quite significant incident angle dependence of R/T ratio).

Current alignment scheme:
Current alignment scheme I figured out is the following.
 - Check Y green. If it is transmitted at good spot on GTRY camera, Yarm is OK. If not, tweak ITMY/ETMY. alignment.
 - Mis-align AS4, align TT1, TT2, LO1 to have DCPD_A_OUT of ~130 and DCPD_B_OUT of ~125.
 - Align PR3, PR2 to maximize TRY_OUT to ~1.05.
 - Tweak ITMY/ETMY if the beam spot on them are not good.
 - Align BS, ITMX to have good MICH fringe and TRX_OUT to ~1.1.
 - Tweak ITMX/ETMX if the beam spot on them are not good.
 - Misalign ETMY, ETMX, ITMY to have LO-ITMX fringe in BHD DCPDs, and align AS beam with SR2 and AS4 differentially, with ratio of AS4/SR2=3.6.

DC PD values in various configurations:
Both arms locked with POX/POY, MICH free, PRM/SRM misaligned

                          Mean     Max      Min
C1:IOO-MC_TRANS_SUM :     14088.57 13947.52 14167.04
C1:LSC-ASDC_OUT16 :           0.16    -0.02     0.34
C1:LSC-POPDC_OUT16 :        369.34   -74.88   854.34
C1:LSC-REFLDC_OUT16 :         0.03    -0.00     0.06
C1:LSC-TRY_OUT16 :            1.00     0.95     1.04
C1:LSC-TRX_OUT16 :            1.07     1.04     1.08

Only LO beam to BHD DCPDs
                          Mean     Max      Min
C1:IOO-MC_TRANS_SUM :     14121.32 14057.71 14159.38
C1:HPC-DCPD_A_OUT16 :       129.80   128.37   130.68 (Consistent with, 40m/17046. Power as expected within 20%. Squashed shape)
C1:HPC-DCPD_B_OUT16 :       123.42   121.92   124.48

ITMX single bounce (ITMY, ETMX, ETMY, PRM, SRM, LO misalgined)
                          Mean     Max      Min
C1:IOO-MC_TRANS_SUM :     14105.13 14000.89 14171.91
C1:HPC-DCPD_A_OUT16 :        92.54    91.45    93.30 (Consistent with 40m/17040, Power as expected within 40%. Clipped to the left in camera)
C1:HPC-DCPD_B_OUT16 :       137.70   136.55   138.53 (Note that DCPD_A/B ratio is different from LO, due to BHD BS R/T unbalance; 40m/17044)
C1:LSC-ASDC_OUT16 :           0.10     0.09     0.10 (Power as expected 40m/16952. Clipped to the right in camera)
C1:LSC-POPDC_OUT16 :        309.19   288.93   327.10 (Power as expected within 30% 40m/17042.)
C1:LSC-REFLDC_OUT16 :         0.02     0.01     0.02

ITMY single bounce (ITMX, ETMX, ETMY, PRM, SRM, LO misalgined)
                          Mean     Max      Min
C1:IOO-MC_TRANS_SUM :     14112.09 14025.37 14154.51
C1:HPC-DCPD_A_OUT16 :        92.58    92.01    93.26
C1:HPC-DCPD_B_OUT16 :       137.68   136.81   138.27
C1:LSC-ASDC_OUT16 :           0.10     0.09     0.10
C1:LSC-POPDC_OUT16 :        308.48   290.49   319.73
C1:LSC-REFLDC_OUT16 :         0.02     0.01     0.02

MICH fringe only (ETMX, ETMY, PRM, SRM, LO misalgined)
                          Mean     Max      Min
C1:IOO-MC_TRANS_SUM :     14090.34 13979.15 14143.86
C1:HPC-DCPD_A_OUT16 :       325.60    91.92   714.57
C1:HPC-DCPD_B_OUT16 :       400.27    18.37   762.57
C1:LSC-ASDC_OUT16 :           0.19    -0.05     0.41
C1:LSC-POPDC_OUT16 :        595.66  -119.21  1334.11
C1:LSC-REFLDC_OUT16 :         0.03    -0.01     0.07

LO-ITMX fringe only (ITMY, ETMX, ETMY, PRM, SRM misalgined)
                          Mean     Max      Min
C1:IOO-MC_TRANS_SUM :     14062.58 13968.05 14113.67
C1:HPC-DCPD_A_OUT16 :       224.31    89.57   371.66
C1:HPC-DCPD_B_OUT16 :       259.74    85.37   421.86


Next:
 - Measure contrast (40m/17020) and estimate mode-matching of LO-AS again (40m/17041)
 - Now that we have better LO-AS fringe, lock LO phase in MICH (40m/17037)
 - Now that Dolphin issue was fixed, try double-demodulation to lock LO phase

Attachment 1: Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png  36 kB  Uploaded Wed Aug 3 17:19:11 2022  | Hide | Hide all
Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
ELOG V3.1.3-